OpenStudy (anonymous):

equation -5y-9=-(y-1)

OpenStudy (anonymous):

you want to solve for y? looking at the RHS of equation, you can apply Distributive Property for the Negative Sign\[-5y-9=-y+1\]Add +y+9 on both side \[-5y+y\bcancel{-9}\bcancel{+9}=\bcancel{-y}\bcancel{+y}+1+9\]\[-4y=10\]Multiply \(-1/4\) both side\[\bcancel{-\left(\frac{1}{4}\right)}\bcancel{-4}y=10\left(-\frac{1}{4}\right)\]or\[y=-\frac{10}{4}\]